OpenStudy (anonymous):

One of two urns is chosen at random with one as likely to be chosen as the other. Then a ball is drawn from the chosen urn. Urn 1 contains 1 white and 3 red balls, and Urn 2 has 2 whites, and 2 red balls. If a white ball is drawn what is the probability that it came from urn 2?

OpenStudy (anonymous):

please help me i dont understand

OpenStudy (anonymous):

// conditional probability P(urn 2|white ball) = P(from urn 2 AND white ball)/P(white ball)

OpenStudy (anonymous):

@Anita505 you got this?

OpenStudy (anonymous):

no can u please expand

OpenStudy (anonymous):

put \(A\) as the event you pick from urn 2 \(B\) as the event you get a white ball. you want \[P(A|B)\] so you understand that notation ?

OpenStudy (anonymous):

*DO you understand that notation is what i meant

OpenStudy (anonymous):

how do i plug in the numbers into that notation?

OpenStudy (anonymous):

you do not. you have to know that \[P(A|B)=\frac{P(A\cap B)}{P(B)}\]

OpenStudy (anonymous):

ok then how do i find the answer from that?

OpenStudy (anonymous):

so you need both \(P(A\cap B)\) and \(P(B)\)

OpenStudy (anonymous):

is the answer 2?

OpenStudy (anonymous):

you can't have a probability greater than 1

OpenStudy (anonymous):

hmm i see that you are somewhat confused 2 is not the answer to any probability question

OpenStudy (anonymous):

the easy part is computing \(P(A\cap B)\) finding \(P(B)\) is a bit harder

OpenStudy (anonymous):

okay so the answer is 4 for P(AnB)

OpenStudy (anonymous):

P(A AND B) = 2/4 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but it will be 1/2 if u simplify it

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

not to butt in again, but i think \[P(A\cap B)=\frac{1}{2}\times \frac{1}{2}=\frac{1}{4}\]

OpenStudy (anonymous):

that is my answer 1/4?

OpenStudy (anonymous):

you have to pick from urn 2 AND get a white ball

OpenStudy (anonymous):

no, that is the numerator of \[\frac{P(A\cap B)}{P(B)}\]

OpenStudy (anonymous):

is the first 1/2 from the fact that you can choose either urn

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\(A\cap B\) means you pick from urn 2 AND get white ball, so both things have to happen

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the answer is 1/4 then

OpenStudy (anonymous):

now comes the harder part, finding \(P(B)\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how do i find P(B) ?

OpenStudy (anonymous):

how can you pick a white ball? 1) you pick from urn 1 AND get a white ball OR 2) you pick from urn 2 AND get a white ball

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

P(from urn 1 AND white ball) + P(from urn 2 AND white ball)

OpenStudy (anonymous):

the probability you pick from urn 1 and get a white ball is \[\frac{1}{2}\times \frac{1}{4}=\frac{1}{8}\]the probability you pick from urn 2 and get a white ball we just computed, it is \(\frac{1}{4}\) add them up and get \[\frac{1}{4}+\frac{1}{8}=\frac{3}{8}\]

OpenStudy (anonymous):

your final answer is therefore \[\frac{\frac{1}{4}}{\frac{3}{8}}=\frac{1}{4}\times \frac{8}{3}=\frac{2}{3}\]

OpenStudy (anonymous):

2/3?